A train leaves Danville Union and travels north at a speed of 75 km/h. Two hours later, an express train leaves on a parallel track and travels north at 125 km/h. How far from the station will they meet?
Answer by checkley77(12844)

D=RT
D=75T
D=125(T-2)
75T=125(T-2)
75T=125T-250
75T-125T=-250
-50T=-250
T=-250/-50
T=5 HOURS IS THE TIME UNTIL THEY MEET.
PROOF:
75*5=125(5-2)
375=125*3
375=375